Circular No. 6957

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998cs IN UGC 10432 T. Puckett, Mountain Town, GA, reports his discovery of an appar-

ent supernova (mag 17.5) on an unfiltered CCD frame taken with the Puckett Observatory 0.30-m automated supernova patrol telescope on June 29.25 UT. The candidate is located at α = 16^{h}30^{m}39^{s}.22, δ = +41^{o}12′54′′.4 (equinox 2000.0), which is 0′′.2 east and 16′′.0 north of the nucleus of UGC 10432. SN 1998cs was also present on CCD frames taken on June 30.14 at mag 17.0. The object was not present on previous CCD frames taken on June 2.17 (limiting mag 18.75). The object does not appear on the Palomar Sky Survey scans.

SUPERNOVA 1998ct IN UGC 10062 M. Schwartz, Cottage Grove, OR, reports his discovery of an apparent

supernova (mag 16.1) on an unfiltered CCD frame taken with the 0.35m Tenagra I automated supernova patrol telescope on June 29.30 UT. The candidate is located at α = 15^{h}51^{m}11^{s}.67, δ = +21^{o}56′36′′.0 (equinox 2000.0), which is 5′′ west and 2′′ north of the center of UGC 10062. The object is also present with unchanged position on a frame taken on June 30.26. The limiting magnitude of the exposures was ∼ 19. SN 1998ct does not appear on either the first or the second Palomar Sky Survey scans, and it is also absent from the USNO A1.0 catalogue.

XTE J2123–058 J. A. Tomsick, J. P. Halpern, and K. M. Leighly, Columbia University;

and E. Perlman, Space Telescope Science Institute, report the identification of the optical counterpart within the ASM (cf. IAUC 6955) and PCA error regions, based on photometry with the MDM Observatory 1.3-m telescope beginning on June 30.31 UT: “The star is at α = 21^{h}23^{m}14^{s}.54, δ = −5^{o}47′52′′.9 (equinox 2000.0; position uncertainty 1′′.0). Magnitudes measured on June 30.44 are U = 16.40, B = 17.28, V = 17.30, R = 17.24, I = 17.22, with preliminary errors of 0.05 mag. This star is only marginally visible on a digitized U.K. Schmidt plate. A low-resolution spectrum obtained on the Kitt Peak 2.1-m telescope shows weak absorption lines of Hβ (equivalent width 0.12 nm), Hγ (EW = 0.41 nm), and He I 447-nm (EW = 0.4 nm). He II 468-nm emission of EW = 0.40 nm is also seen. These properties are typical of x-ray transients in outburst.”

1998 June 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green